OpenStudy (anonymous):

A rocket engine, 2.39 cross product 10^−5 watts/cm^2

OpenStudy (anonymous):

dB=10log[I/(10^-12)] Not much to do here, just plug in the value for I (intensity). First we have to convert from W/cm^2 to W/m^2. Just multiply it by 100^2, because there are 100^2 cm^2/m^2 dB=10log[(.266)/(10^-12)] =10log(2.66e11) =114 dB

OpenStudy (anonymous):

hold on...I messed up with something

OpenStudy (anonymous):

dB=10log(I/I0) =10log(2.46∗10−1/10−12) =10log(2.46∗1011)=113.91dB−−−−−−−−−ans

OpenStudy (anonymous):

A rocket engine, 2.39 cross 10−5watts/cm2

OpenStudy (anonymous):

Hold on..even I'm confused now...I'm going to ask @Catlover5925 for help

OpenStudy (anonymous):

@roneal I asked her for help..hold on

OpenStudy (catlover5925):

idk, these are hard questions i dont understand what you are doing here